# Proximal Interacting Particle Langevin Algorithms

### v1

@@ A.4.1 Error bound for proximal gradient Langevin algorithm
$2\gamma\operatorname*{KL}(\mathcal{L}({Y_{n}})\;|\;\pi_{\lambda})\leq$ $(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})-\frac{\gamma}{\lambda}W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})$
$-\left(1-\frac{\gamma}{\lambda}\right)W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi_{\lambda})+\gamma(2\gamma L_{g_{1}}d+\lambda C).$
@@
Since $g_{1}+g_{2}^{\lambda}$ is convex by Durmus et al. [2019, Lemma 1] the following holds $\pi_{\lambda}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathcal{H}(\pi_{\lambda})<\infty$, $\mathcal{E}_{g_{1}+g_{2}^{\lambda}}(\pi_{\lambda})<\infty$ and for all $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$ satisfying $\mathcal{E}_{g_{1}+g_{2}^{\lambda}}(\pi)<\infty$,
$\operatorname*{KL}(\pi\;|\;\pi_{\lambda})=\mathcal{E}_{g_{1}+g_{2}^{\lambda}}(\pi)+\mathcal{H}(\pi)-(\mathcal{E}_{g_{1}+g_{2}^{\lambda}}(\pi_{\lambda})+\mathcal{H}(\pi_{\lambda}))=\mathcal{F}(\pi)-\mathcal{F}(\pi_{\lambda}).$
We can further decompose $\mathcal{E}_{g_{1}+g_{2}^{\lambda}}(\pi)=\mathcal{E}_{g_{1}}(\pi)+\mathcal{E}_{g_{2}^{\lambda}}(\pi)$. Using Durmus et al. [2019, Lemma 5] we have that the negative entropy satisfies the following inequality
$2\gamma\big{[}\mathcal{H}(\mathcal{L}({Y_{n}}))-\mathcal{H}(\pi_{\lambda})\big{]}\leq W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda})-W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi_{\lambda}).$ (33)
@@
Note that $Y_{n}-Z_{n}=\sqrt{2\gamma}\xi_{n}$ is independent of $Z_{n}$, $\mathbb{E}[Y_{n}-Z_{n}]=0$ and $\mathbb{E}[\|Y_{n}-Z_{n}\|^{2}]=2\gamma d$, where $d$ is the dimension of the random variable $\xi_{n}$. Therefore, taking expectations in the previous inequality we get
@@
$2\gamma\big{[}\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))-\mathcal{E}_{g_{1}}(\pi_{\lambda})\big{]}$ $\leq(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})-W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda}),$ (35)
$2\gamma\big{[}\mathcal{E}_{g_{2}^{\lambda}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}^{\lambda}}(\pi_{\lambda})\big{]}$ $\leq\frac{\gamma}{\lambda}W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi_{\lambda})-\frac{\gamma}{\lambda}W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})+\gamma\lambda C.$ (36)
Summing up (33)-(36) and using that $\operatorname*{KL}(\mathcal{L}({Y_{n}})\;|\;\pi_{\lambda})=\mathcal{F}(\mathcal{L}({Y_{n}}))-\mathcal{F}(\pi_{\lambda})$ we have the desired result.
@@
$W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({X_{0}}),\pi_{\lambda})+\frac{\lambda(2\gamma L_{g_{1}}d+\lambda C)}{1-\lambda(1-\mu\gamma)/\gamma}.$
@@
$W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})\leq\frac{\lambda(1-\gamma\mu)}{\gamma}W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})+\lambda(2\gamma L_{g_{1}}d+\lambda C).$
@@
$W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})$ $\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({X_{0}}),\pi_{\lambda})+\lambda(2\gamma L_{g_{1}}d+\lambda C)\sum_{i=0}^{n-1}\frac{\lambda^{i}(1-\gamma\mu)^{i}}{\gamma^{i}}$
$\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({X_{0}}),\pi_{\lambda})+\frac{\lambda(2\gamma L_{g_{1}}d+\lambda C)}{1-\lambda(1-\gamma\mu)/\gamma},$
@@
Remark A.1.
In Theorem A.3 and Corollary A.4, we did not use the convexity of $g_{2}$ but instead that of $g_{2}^{\lambda}$. Therefore, if we consider the following formulation of the algorithm
$X_{n+1/2}$ $=X_{n}-\gamma\nabla g_{1}(X_{n})+\sqrt{2\gamma}\;\xi_{n}$
$X_{n+1}$ $=X_{n+1/2}-\lambda\nabla g_{2}^{\lambda}(X_{n+1/2}),$
it suffices to assume that $g_{2}^{\lambda}$ is convex without the need to make any assumptions about $g_{2}$$\lozenge$
@@ A.4.2 Convergence and discretisation bounds
$V_{n+1/2}$ $=V_{n}-\gamma\nabla g_{1}(V_{n})+\sqrt{\frac{2\gamma}{N}}\;\xi_{n},$
$V_{n+1}$ $=V_{n+1/2}-\lambda\nabla g_{2}^{\lambda}(V_{n+1/2}).$ (37)
@@
$W_{2}^{2}(\mathcal{L}({V_{n}}),\pi_{\lambda}^{N})\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({V_{0}}),\pi_{\lambda}^{N})+\frac{\lambda(2\gamma L_{g_{1}}d+\lambda C)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}.$
@@
where $\tilde{\gamma}=\gamma/N$ and $\tilde{\lambda}=\lambda/N$. Since $\mathbb{E}[\|\nabla g_{2}^{\lambda}(V_{n+1/2})\|^{2}]\leq C$ for all $n\in\mathbb{N}$ by D3, it follows that $\mathbb{E}[\|\nabla\tilde{g}_{2}^{\tilde{\lambda}}(V_{n+1/2})\|^{2}]\leq NC$. Therefore, taking $\tilde{\gamma}\leq 1/(NL_{g_{1}})$ which is equivalent to $\gamma\leq 1/L_{g_{1}}$, and applying Corollary A.4 the result follows. ∎
@@
$\nabla G_{1}(z_{\theta},z_{1},\dots,z_{N})=\Big{(}N^{-1}\sum_{i=1}^{N}\nabla_{\theta}g_{1}(z_{\theta},\sqrt{N}z_{i}),N^{-1/2}\nabla_{z_{1}}g_{1}(z_{\theta},\sqrt{N}z_{1}),\dots,N^{-1/2}\nabla_{z_{N}}g_{1}(z_{\theta},\sqrt{N}z_{N})\Big{)}^{\intercal}$
@@
Let $Z_{n}=(\theta_{n}^{N},N^{-1/2}X_{n}^{1,N},\dots,N^{-1/2}X_{n}^{N,N})$ and $\pi_{\lambda}^{N}\propto\exp(-N(G_{1}+G_{2}^{\lambda}))$. Suppose that A1, A4 and C1 hold true and assume $\gamma\leq 1/L_{g_{1}}$ and $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$. Then,
$W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda}^{N})\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({Z_{0}}),\pi_{\lambda}^{N})+\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda C)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}.$
@@
Note that if C1 holds then $\mathbb{E}[\|\nabla G_{2}^{\lambda}(z)\|^{2}]\leq C$. To see this consider
$\|\nabla G_{2}^{\lambda}(z)\|^{2}=\Big{\|}\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}g_{2}^{\lambda}(z_{\theta},\sqrt{N}z_{i})\Big{\|}^{2}+\frac{1}{N}\sum_{i=1}^{N}\big{\|}\nabla_{x}g_{2}^{\lambda}(z_{\theta},\sqrt{N}z_{i})\big{\|}^{2}.$
Using the fact that $(N^{-1}\sum_{i}a_{i})^{2}\leq N^{-1}\sum_{i}a_{i}^{2}$, we get
$\|\nabla G_{2}^{\lambda}(z)\|^{2}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\left(\big{\|}\nabla_{\theta}g_{2}^{\lambda}(z_{\theta},\sqrt{N}z_{i})\big{\|}^{2}+\big{\|}\nabla_{x}g_{2}^{\lambda}(z_{\theta},\sqrt{N}z_{i})\big{\|}^{2}\right)=\frac{1}{N}\sum_{i=1}^{N}\big{\|}\nabla g_{2}^{\lambda}(z_{\theta},\sqrt{N}z_{i})\big{\|}^{2}.$
Finally, taking expectations in both sides we obtain
$\mathbb{E}[\|\nabla G_{2}^{\lambda}(Z_{n+1/2})\|^{2}]\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\big{[}\big{\|}\nabla g_{2}^{\lambda}(\theta_{n+1/2}^{N},X^{i,N}_{n+1/2})\big{\|}^{2}\big{]}\leq C.$
In addition, observe that A1, A4 imply that $G_{1}$ and $G_{2}^{\lambda}$ are convex since $g_{1}$ and $g_{2}^{\lambda}$ are convex, and $G_{1}$ is also $\mu$-strongly convex and $L_{g_{1}}$-gradient Lipschitz.
The proof then follows from Corollary A.5. ∎
@@
$W_{2}(\mathcal{L}({\theta^{N}_{n}}),\delta_{\bar{\theta}_{\star}})\leq$ $\frac{\lambda}{\mu}\left(\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\right)+\sqrt{\frac{2d_{\theta}}{N\mu}}+\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi_{\lambda}^{N})$
$+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda C)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2}+\mathcal{O}(\lambda^{2})$
@@
Note that $\pi_{\lambda}^{N}(z)\propto\exp(-N(G_{1}(z)+G_{2}^{\lambda}(z)))=\exp(-\sum_{i}U^{\lambda}(z_{\theta},\sqrt{N}z_{i}))$. By Corollary A.6 it follows that
$W_{2}(\mathcal{L}({\theta_{n}^{N}}),\pi_{\lambda,\Theta}^{N})\leq W_{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda}^{N})\leq\sqrt{\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({Z_{0}^{N}}),\pi_{\lambda}^{N})+\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda C)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}}.$
@@
$W_{2}(\mathcal{L}({\theta_{n}^{N}}),\pi_{\lambda,\Theta}^{N})\leq\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi_{\lambda}^{N})+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda C)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2}.$
The proof then follows from (31), Proposition A.1 and the above. ∎

### v2

@@ A.4.1 Error bound for proximal gradient Langevin algorithm
$2\gamma\operatorname*{KL}(\mathcal{L}({Y_{n}})\;|\;\pi)\leq$ $(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)-\frac{\gamma}{\lambda}W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi)$
$-\left(1-\frac{\gamma}{\lambda}\right)W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi)+\gamma(2\gamma L_{g_{1}}d+\lambda C).$
@@
Since $g_{1}+g_{2}$ is convex by assumption the following holds $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathcal{H}(\pi)<\infty$, $\mathcal{E}_{g_{1}+g_{2}}(\pi)<\infty$ and for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ satisfying $\mathcal{E}_{g_{1}+g_{2}}(\mu)<\infty$,
$\operatorname*{KL}(\mu\;|\;\pi)=\mathcal{E}_{g_{1}+g_{2}}(\mu)+\mathcal{H}(\mu)-(\mathcal{E}_{g_{1}+g_{2}}(\pi)+\mathcal{H}(\pi))=\mathcal{F}(\mu)-\mathcal{F}(\pi).$
We can further decompose $\mathcal{E}_{g_{1}+g_{2}}(\mu)=\mathcal{E}_{g_{1}}(\mu)+\mathcal{E}_{g_{2}}(\mu)$. Using Durmus et al. [2019, Lemma 5] we have that the negative entropy satisfies the following inequality
$2\gamma\big{[}\mathcal{H}(\mathcal{L}({Y_{n}}))-\mathcal{H}(\pi)\big{]}\leq W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi)-W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi).$ (35)
@@
Note that $Y_{n}-Z_{n}=\sqrt{2\gamma}\xi_{n}$ is independent of $Z_{n}$, $\mathbb{E}[Y_{n}-Z_{n}]=0$ and $\mathbb{E}[\|Y_{n}-Z_{n}\|^{2}]=2\gamma d$, where $d$ is the dimension of the standard Gaussian random variable $\xi_{n}$. Therefore, taking expectations in the previous inequality we get
@@
$2\gamma\big{[}\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))-\mathcal{E}_{g_{1}}(\pi)\big{]}$ $\leq(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)-W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi),$ (37)
$2\gamma\big{[}\mathcal{E}_{g_{2}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}}(\pi)\big{]}$ $\leq\frac{\gamma}{\lambda}W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi)-\frac{\gamma}{\lambda}W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi)+\gamma\lambda C.$ (38)
Summing up (35)-(38) and using that $\operatorname*{KL}(\mathcal{L}({Y_{n}})\;|\;\pi)=\mathcal{F}(\mathcal{L}({Y_{n}}))-\mathcal{F}(\pi)$ we have the desired result.
@@
$W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({X_{0}}),\pi)+\frac{\lambda(2\gamma L_{g_{1}}d+\lambda C)}{1-\lambda(1-\mu\gamma)/\gamma}.$
@@
$W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi)\leq\frac{\lambda(1-\gamma\mu)}{\gamma}W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)+\lambda(2\gamma L_{g_{1}}d+\lambda C).$
@@
$W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)$ $\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({X_{0}}),\pi)+\lambda(2\gamma L_{g_{1}}d+\lambda C)\sum_{i=0}^{n-1}\frac{\lambda^{i}(1-\gamma\mu)^{i}}{\gamma^{i}}$
$\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({X_{0}}),\pi)+\frac{\lambda(2\gamma L_{g_{1}}d+\lambda C)}{1-\lambda(1-\gamma\mu)/\gamma},$
@@ A.4.2 Convergence and discretisation bounds
$V_{n+1/2}$ $=V_{n}-\gamma\nabla g_{1}(V_{n})+\sqrt{\frac{2\gamma}{N}}\;\xi_{n},$ (39)
$V_{n+1}$ $=V_{n+1/2}-\lambda\nabla g_{2}^{\lambda}(V_{n+1/2}).$
@@
$W_{2}^{2}(\mathcal{L}({V_{n}}),\pi^{N})\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({V_{0}}),\pi^{N})+\frac{\lambda(2\gamma L_{g_{1}}d+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}.$
@@
where $\tilde{\gamma}=\gamma/N$ and $\tilde{\lambda}=\lambda/N$. Note that the subdifferential set satisfies $\partial\tilde{g}_{2}=N\partial g_{2}$. Therefore, since $\|\nabla^{0}g_{2}(x)\|^{2}\leq C$ for all $x\in\mathbb{R}^{d}$ by D3, it follows that $\|\nabla^{0}\tilde{g}_{2}(x)\|^{2}\leq N^{2}C$. Therefore, taking $\tilde{\gamma}\leq 1/(NL_{g_{1}})$ which is equivalent to $\gamma\leq 1/L_{g_{1}}$, and applying Corollary A.4 the result follows. ∎
@@
$\nabla G_{1}(z_{\theta},z_{1},\dots,z_{N})=$ $\Big{(}N^{-1}\sum_{i=1}^{N}\nabla_{\theta}g_{1}(z_{\theta},\sqrt{N}z_{i}),N^{-1/2}\nabla_{z_{1}}g_{1}(z_{\theta},\sqrt{N}z_{1}),\dots,$
$N^{-1/2}\nabla_{z_{N}}g_{1}(z_{\theta},\sqrt{N}z_{N})\Big{)}^{\intercal}$
@@
Let $Z_{n}=(\theta_{n}^{N},N^{-1/2}X_{n}^{1,N},\dots,N^{-1/2}X_{n}^{N,N})$ and $\pi^{N}\propto\exp(-N(G_{1}+G_{2}))$. Suppose that A1, A4 and C1 hold true and assume $\gamma\leq 1/L_{g_{1}}$ and $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$. Then,
$W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi^{N})\leq\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({Z_{0}}),\pi^{N})+\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}.$
@@
Note that if C1 holds then $\|\nabla^{0}G_{2}(z)\|^{2}\leq C$ for every $z$. To see this note that
$\partial G_{2}(z)=\partial G_{2}(z_{\theta},z_{1},\dots,z_{N})=\frac{1}{N}\sum_{i=1}^{N}\partial g_{2}(z_{\theta},\sqrt{N}z_{i}).$
Therefore, using the fact that $(N^{-1}\sum_{i}a_{i})^{2}\leq N^{-1}\sum_{i}a_{i}^{2}$, we get that the minimal section satisfies
$\|\nabla^{0}G_{2}(z)\|^{2}\leq\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla^{0}g_{2}(z_{\theta},\sqrt{N}z_{i})\right\|^{2}\leq C.$
In addition, observe that A1, A4 imply that $G_{1},G_{2}$ and $G_{2}^{\lambda}$ are convex since $g_{1},g_{2}$ and $g_{2}^{\lambda}$ are convex, and $G_{1}$ is also $\mu$-strongly convex and $L_{g_{1}}$-gradient Lipschitz. The proof then follows from Corollary A.5. ∎
Before proving our final result for $W_{2}(\mathcal{L}(\theta_{n}^{N}),\delta_{\bar{\theta}_{\star}})$, we provide a result adapted from Altschuler and Chewi [2023, Lemma A.8] to our non-differentiable setting that will be useful to bound the first term of (31).
Lemma A.8.
Suppose that the distribution $\pi\propto\exp(-f)$ on $\mathbb{R}^{d}$ is $\alpha$-strongly log-concave, almost everywhere differentiable and that $x^{\star}$ is the minimiser of $f$. Then,
$\mathbb{E}_{X\sim\pi}[\|X-x^{\star}\|^{2}]\leq d/\alpha.$
Proof.
Let $\Omega\subset\mathbb{R}^{d}$ denote the set of differentiable points of $f$, note that when $f$ is convex and differentiable at $x\in\Omega$, then $\partial f(x)=\{\nabla f(x)\}$, that is, its gradient is its only subgradient. Recall also that $f$ is strongly convex, so for every $x,y\in\mathbb{R}^{d}$ we have that
$\langle\partial f(x),x-y\rangle\geq\alpha\|x-y\|^{2}.$
Integration by parts shows that for any smooth function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ of controlled growth, it holds that
$0=\int_{\Omega}\left(\Delta\phi-\langle\nabla f,\nabla\phi\rangle\right)\mathrm{d}\pi=\mathbb{E}_{X\sim\pi}[\Delta\phi-\langle\nabla f,\nabla\phi\rangle].$ (41)
Applying (41) to the function $\phi(x):=\|x-x^{\star}\|^{2}/2$, for which $\nabla\phi(x)=x-x^{\star}$ and $\Delta\phi=d$, together with the strong convexity of $f$, the result follows. ∎
@@
$W_{2}(\mathcal{L}({\theta^{N}_{n}}),\delta_{\bar{\theta}_{\star}})\leq$ $\sqrt{\frac{d_{\theta}}{N\mu}}+\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi^{N})+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2}$
@@
Using a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\Theta}$ is $\mu$-strongly log-concave. Therefore, $\pi_{\Theta}^{N}$ is $N\mu$-strongly log-concave and satisfies all the assumptions of Lemma A.8. So, we have that
$W_{2}(\delta_{\bar{\theta}_{*}},\pi_{\Theta}^{N})^{2}\leq\frac{d_{\theta}}{N\mu}.$
On the other hand, note that $\pi^{N}(z)\propto\exp(-N(G_{1}(z)+G_{2}(z)))=\exp(-\sum_{i}U(z_{\theta},\sqrt{N}z_{i}))$. By Corollary A.6 it follows that
$W_{2}(\mathcal{L}({\theta_{n}^{N}}),\pi_{\Theta}^{N})\leq W_{2}(\mathcal{L}({Z_{n}}),\pi^{N})\leq\sqrt{\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}W_{2}^{2}(\mathcal{L}({Z_{0}^{N}}),\pi^{N})+\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}}.$
@@
$W_{2}(\mathcal{L}({\theta_{n}^{N}}),\pi_{\Theta}^{N})\leq\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi^{N})+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2}.$
The proof then follows from (31) and the above. ∎